OpenStudy (anonymous):

Did I do it right this time? |3-4x|≥7 |-4x+3|≥7 -7≤-4x+3≤7 -10≤-4x≤4 10/4≤x≤-1 final answer: (10/4, -1)

OpenStudy (anonymous):

dang interval wrong again

OpenStudy (anonymous):

gyaaah! darn

OpenStudy (anonymous):

i think i confused you when i said to switch, so maybe it is better if you do not

OpenStudy (anonymous):

ok, so then do I need to switch 10/4 and -1? or is it something else?

OpenStudy (anonymous):

first of all you have an inequality that looks like this: \[|\text{something}|>p\]

OpenStudy (anonymous):

if you see this symbol \[| blah| > p\] you should know that your answer will be TWO intervals, not one

OpenStudy (anonymous):

if you have \[|blah|<p\] one interval \[|blah|> p\] TWO intervals

OpenStudy (anonymous):

less than, one interval. greater than, two intervals. i know i am repeating myself but if you know this in advance it will make your live easier, trust me

OpenStudy (anonymous):

so once you see \[|3-4x|\geq7\] you should right away think "i am going to solve TWO inequalities and get TWO intervals"

OpenStudy (anonymous):

the two inequalities you have so solve are \[3-4x\geq 7\] or \[3-4x\leq -7\] and you have to solve them separately.

OpenStudy (anonymous):

first one is \[3-4x\geq 7\] \[-4x\geq 4\] \[x\leq -1\] note the change in sign when dividing by -4

OpenStudy (anonymous):

second one is \[3-4x\leq -7\] \[-4x\leq -10\] \[x\geq \frac{5}{2}\]

OpenStudy (anonymous):

I forgot to simplify, didn't I?

OpenStudy (anonymous):

and so your answer has two, count them, two intervals, they are \[x\leq -1\] or \[(-\infty,-1)\] or \[x\geq \frac{5}{2}\] or \[(\frac{5}{2},\infty)\]

OpenStudy (anonymous):

yes but you also had one interval when you needed two intervals

OpenStudy (anonymous):

that to what is the purpose of the infinite?

OpenStudy (anonymous):

you can write \[(-\infty,-1)\cup (\frac{5}{2},\infty)\] if you like

OpenStudy (anonymous):

how do you write \[x\leq 1\] in interval notation?

OpenStudy (anonymous):

it means x can be any number that is less than one. anything.

OpenStudy (anonymous):

Ohhh ok. That makes sense. Thank you again!

OpenStudy (anonymous):

yw. and don't forget please: absolute value is LESS THAN gives ONE INTERVAL absolute value GREATER THAN gives TWO INTERVALS

OpenStudy (anonymous):

Wait, what is the U symbol? and I wont!

OpenStudy (anonymous):

that means "union" so it could be in one set OR the other

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

you could just write the word OR instead

OpenStudy (anonymous):

try another one carefully and see if you get it right. your algebra is correct, intervals are a problem

OpenStudy (anonymous):

shouldnt we be using "[", "]" instead of "(",")" (on this side of the numbers in each interval - not for infinity) since we have "<=" and ">=" here ?

OpenStudy (anonymous):

i.e.: \[(-\infty,-1]\cup[\frac{5}{2},\infty)\]

OpenStudy (anonymous):

yes of course you are right. i totally ignored that party. sorry. so much to think about...